OpenStudy (anonymous):

The lawn in front of Kermit Middle School is in the shape of a rectangle 30 yd long and 16 yd wide. How much shorter if you walk diagonally across the lawn rather than along two sides of it?

OpenStudy (anonymous):

Use the Pythagorean Theorem: a^2 + b^2 = c^2

OpenStudy (anonymous):

yeahh i got that part....but its asking for "HOW MUCH SHORTER IF U WALK DIAGONALLY.....?"

OpenStudy (anonymous):

when you slice a rectangle diagonally, you get a triangle. when you find the 3 sides, you can compare the 3 numbers.

OpenStudy (anonymous):

what? i got 34 yds for the diagonal...i dont understand how to find how much shorter it is to walk diagonally.

OpenStudy (anonymous):

hey can someone help me out?

OpenStudy (anonymous):

add sides and subtract diagonal ...i think

OpenStudy (anonymous):

i tried..it doesn't work.

OpenStudy (anonymous):

is this multiple choice? if so, show me the answers

OpenStudy (anonymous):

no, its not multiple choice.

OpenStudy (anonymous):

Try X = 30squared + 16squared = 900 + 256 =556 square root it to get your answer

OpenStudy (anonymous):

23.58 (2dp)

OpenStudy (anonymous):

Hope that helped =)

OpenStudy (turingtest):

The answer is 12 let me know if you want the work shown

OpenStudy (anonymous):

so i did explain this correctly... 46 (total) - 34 (diagonal) = 12

OpenStudy (turingtest):

yeah, he just didn't do it right probably